(2x+3)/4 + (x-5) / 3 = 3/2
I have done some of this but i have gotten stuck after this step:
2x+3/4 x 3 = 6x + 9 / 12
x-5 / 3 x 4 = 4x - 20 / 2
6x + 4x = 10x
+9 - 20 = -11
so i now have,
10x - 11 = 3/2, please help me work this out, can you please show me how to finish it using an easier method than the calculator,
Many thanks!

Answer by Cromlix(4381)

(2x+3)/4 + (x-5) / 3 = 3/2
Lowest Common Denominator = 12
3(2x + 3) + 4(x - 5) = 18
I have divided all the different
denominators into 12 and multiplied
the numerators by the answer.
Multiply out:
6x + 9 + 4x - 20 = 18
Collect like terms
6x + 4x = 18 - 9 + 20
10x = 29
x = 29/10 OR 2 9/10
Hope this helps.
:-)